The Neal Company wants to estimate next year's return on equity (ROE) under different financial leverage ratios. Neal's total capital is $16 million, it currently uses only common equity, it has no future plans to use preferred stock in its capital structure, and its federal-plus-state tax rate is 40%. The CFO has estimated next year's EBIT for three possible states of the world: $4.2 million with a 0.2 probability, $2 million with a 0.5 probability, and $0.7 million with a 0.3 probability. Calculate Neal's expected ROE, standard deviation, and coefficient of variation for each of the following debt-to-capital ratios. Do not round intermediate calculations. Round your answers to two decimal places at the end of the calculations.

Debt/Capital ratio is 0.

RÔE = ___%
σ = ___%
CV =

Debt/Capital ratio is 10%, interest rate is 9%.

RÔE = ___%
σ = ___%
CV =

Debt/Capital ratio is 50%, interest rate is 11%.

RÔE = ___%
σ = ___%
CV =

Debt/Capital ratio is 60%, interest rate is 14%.

RÔE = ___%
σ = ___%
CV =
Answer #1

Ans.

Debt/Capital ratio is 0
Probability EBIT Equity Debt Interest Net income= (EBIT-interest)*(1-tax) ROE (Net Income / Equity)
0.2 $4,200,000 $16,000,000 0 0 $2,520,000 ($4,200,000-$0)*(1-0.40) 15.75%
0.5 $2,000,000 $16,000,000 0 0 $1,200,000($2,000,000-$0)*(1-0.40) 7.50%
0.3 $700,000 $16,000,000 0 0 $420,000 ($700,000-$0)*(1-0.40) 2.63%

Expected ROE = 0.2*15.75% + 0.5*7.50% +0.3*2.63%

Expected ROE = 3.15% + 3.75% + 0.79% = 7.69%

Standard Deviation = Square root (0.2*(15.75%-7.69%)2 + 0.5*(7.5%-7.69%)2 +0.3*(2.63%-7.69%)2)

Standard Deviation = 4.55%

CV = Standard Deviation/Expected ROE = 4.55% / 7.69%

CV = 0.59

Debt/Capital ratio is 10%, interest rate is 9%
Probability EBIT Equity Debt Interest Net income= (EBIT-interest)*(1-tax) ROE (Net Income / Equity)
0.2 $4,200,000 $14,400,000 $1,600,000 $144,000 $2,433,600 ($4,200,000-$144,000)*(1-0.40) 16.90%
0.5 $2,000,000 $14,400,000 $1,600,000 $144,000 $1,113,600($2,000,000-$144,000)*(1-0.40) 7.73%
0.3 $700,000 $14,400,000 $1,600,000 $144,000 $333,600 ($700,000-$144,000)*(1-0.40) 2.32%

Expected ROE = 0.2*16.9% + 0.5*7.73% +0.3*2.32%

Expected ROE = 3.38% + 3.87% + 0.70% = 7.94%

Standard Deviation = Square root (0.2*(16.9%-7.94%)2 + 0.5*(7.73%-7.94%)2 +0.3*(2.32%-7.94%)2)

Standard Deviation = 5.05%

CV = 5.05% / 7.94%

CV = 0.64

Debt/Capital ratio is 50%, interest rate is 11%
Probability EBIT Equity Debt Interest Net income= (EBIT-interest)*(1-tax) ROE (Net Income / Equity)
0.2 $4,200,000 $8,000,000 $8,000,000 $880,000 $1,992,000 ($4,200,000-$880,000)*(1-0.40) 24.90%
0.5 $2,000,000 $8,000,000 $8,000,000 $880,000 $672,000($2,000,000-$880,000)*(1-0.40) 8.40%
0.3 $700,000 $8,000,000 $8,000,000 $880,000 -$108,000 ($700,000-$880,000)*(1-0.40) -1.35%

Expected ROE = 0.2*24.9% + 0.5*8.40% +0.3* (- 1.35)%

Expected ROE = 4.98% + 4.2 % - 0.41 % = 8.78%

Standard Deviation = Square root (0.2*(24.9%-8.78%)2 + 0.5*(8.40% - 8.78%)2 +0.3*(-1.35%-8.78%)2)

Standard Deviation = 9.10%

CV = 9.10% / 8.78%

CV = 1.04

Debt/Capital ratio is 60%, interest rate is 14%
Probability EBIT Equity Debt Interest Net income= (EBIT-interest)*(1-tax) ROE (Net Income / Equity)
0.2 $4,200,000 $6,400,000 $9,600,000 $1,344,000 $1,713,600 ($4,200,000-$1,344,000)*(1-0.40) 26.78%
0.5 $2,000,000 $6,400,000 $9,600,000 $1,344,000 $393,600($2,000,000-$1,344,000)*(1-0.40) 6.15%
0.3 $700,000 $6,400,000 $9,600,000 $1,344,000 $386,400 ($700,000-$1,344,000)*(1-0.40) -6.04%

Expected ROE = 0.2*26.78% + 0.5*6.15% +0.3*(-6.04)%

Expected ROE = 5.36% + 3.08 % - 1.81 % = 6.62%

Standard Deviation = Square root (0.2*(26.78%-6.62%)2 + 0.5*(6.15%-6.62%)2 +0.3*(-6.04% - 6.62 %)2)

Standard Deviation = 11.38%

CV = 11.38% / 6.62%

CV = 1.72